OpenStudy (anonymous):

Let F(x)=int_{0}^{x}(e^{{3t})^{4}} Find the MacLaurin polynomial of degree 5 for F(x)

OpenStudy (anonymous):

please help

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

i dont exactly know how to put in the answer

OpenStudy (zarkon):

what do you mean 'put in the answer'?

OpenStudy (anonymous):

ok please solve it and i can see your answer then compare to mine

OpenStudy (zarkon):

why don't you give me yours and I'll tell you if you are correct

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

629856 e^(81 t^4) t^3 (35+10530 t^4+524880 t^8+5668704 t^12)

OpenStudy (anonymous):

thats e^(−3t)^4 to the 5th degree

OpenStudy (zarkon):

why do you have t's

OpenStudy (zarkon):

F is a function of x

OpenStudy (anonymous):

well i do not know how to integrate e^(−3t)^4

OpenStudy (zarkon):

you don't need to

OpenStudy (zarkon):

just for clarification is it \[\int\limits_{0}^{x}e^{(3t)^4}dt\] or \[\int\limits_{0}^{x}\left(e^{3t}\right)^4dt\]

OpenStudy (anonymous):

the first one

OpenStudy (zarkon):

ok..I'll get you started... \[F(0)=\int\limits_{0}^{0}e^{(3t)^4}dt=0\] \[\frac{d}{dx}F(x)=\frac{d}{dx}\int\limits_{0}^{x}e^{(3t)^4}dt=e^{(3x)^4}\] \[F'(0)=1\]

OpenStudy (anonymous):

972e^(81x^4) (x^2+108x^6)

OpenStudy (anonymous):

is that the second derivative ?

OpenStudy (anonymous):

?

OpenStudy (zarkon):

\[324x^3e^{81x^4}\]

OpenStudy (anonymous):

do you mind explaining that, because i double checked with wolf ram alpha and i got my answer

OpenStudy (anonymous):

actually your answer is the first derivative. Isnt it ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so is my answer 1944e^(81x^4)(1+(2754x^4)+(314928x^8)+5668704x^12)

OpenStudy (anonymous):

or f(0) of that which is 1944

OpenStudy (anonymous):

because according to my program, none of those answers are right

OpenStudy (anonymous):

?

OpenStudy (zarkon):

972e^(81x^4) (x^2+108x^6) is the 3rd derivative not the 2nd